# StoneDetector: Conventional and versatile code clone detection for Java

### v1

@@ 3 Architecture and Implementation
Apparently, as can be seen by the above discussion, the StoneDetector platform provides a rich set of configuration parameters that can be used for optimizing clone detection and demonstrates the platform’s versatility. Among the most critical parameters are the threshold value $\tau$ defining the required similarity of code clones, the minimum size $l$ of code fragments acceptable as code clones, the string metric $m$ used for path matching, and the hash function h used for description set encoding (cf. also Table 1). For an in-depth description of the platform’s configuration parameters, we refer the reader to the supplementary information available online [^4] https://stonedetector.fmi.uni-jena.de/.
@@
String metric $\delta$ LCS LCS
@@
Hashing algorithm $h$ none MD5
@@
Table 1: Main configuration parameters in the StoneDetector platform
3.1 Extensibility of the StoneDetector platform
@@
The use of fingerprints to represent description sets makes clone detection in the StoneDetector platform fundamentally independent of the selected programming language and possible program representation. In fact, as is also depicted in Fig. 5, StoneDetector can be divided into a frontend and backend, with the former generating description sets of given code fragments and the latter performing the actual clone matching. The strict separation of frontend and backend allows us to facilitate a readily extensible platform, which can be easily extended for processing other programming languages as well as other internal program representations, such as control flow graphs, abstract syntax trees, or even just token sequences of linear code.
Such an extension is only required to map the new programming language or program representation to fingerprints by using a certain hash function. The clone matching can subsequently be performed using the existing backend of the StoneDetector platform. The specific hash function used for construction is irrelevant to the functionality of the backend. In this way, fingerprints form a general and elegant interface for extending our system to a wide variety of programming languages and program representations.
@@
To demonstrate the platform’s versatility and extensibility, we have added the handling of Java Bytecode as another frontend of the StoneDetector platform [40] (cf. Fig. 5). The new frontend gets a set of class files as input and transforms them to fingerprint sets, which are committed to the existing clone matcher of StoneDetector. Specifically, processing class files is implemented via the Soot framework [41], which first transforms the contained Bytecode into the Bytecode-like Baf or into the register-based Jimple representation. Afterward, dominator trees are generated based upon the Baf or Jimple representation and paths contained in the dominator trees are extracted and encoded into fingerprint sets, as required for clone matching.
@@ 4 Evaluation
In the evaluation, we are interested in finding answers to a number of research questions, in particular about the performance of our code clone detector compared to the current state of the art. We also present results that we believe to provide general insights into code clone detection for Java and to help other researchers when implementing code clone detectors, besides showing the versatility of the StoneDetector platform. Thus, we address the following research questions:
@@
    Research Question 1: How many code clones does our clone detection method and tool StoneDetector detect, i.e., what is its recall compared to the state of the art?
@@
    Research Question 2: How many of StoneDetector’s reported code clones are real code clones, i.e., what is the tool’s precision compared to the state of the art?
@@
    Research Question 3: Are conventional code clone detectors like StoneDetector in general able to find more difficult-to-identify code clones, i.e., code clones with larger syntactical deviation?
@@
    Research Question 4: Does StoneDetector scale to large codebases containing up to multiple million lines of code?
@@
    Research Question 5: How do the used threshold value and the minimal size of code clone influence StoneDetector’s performance?
  • 6.
    Research Question 6: Which role do different string metrics and hashing methods play in a code clone detector like StoneDetector?
  • 7.
    Research Question 7: Can source code clones also be found in Java Bytecode using clone detectors like StoneDetector?
@@
In the comparison of StoneDetector with other code clone detectors for Java, a set of representative state-of-the-art tools has been chosen, including conventional text-based, token-based, tree-based, and data-driven clone detectors, namely, we use the tools:
@@
In the evaluation experiments, we only consider code fragments with a minimal length of 15 lines in the original source code if not otherwise stated. Considering smaller code fragments in general results in a skewed comparison. For instance, in a Java method with only 6 lines of source code, half of the lines are usually just taken by the method’s annotations, signature, and curly braces. However, for the curious reader, we also provide experiments regarding different minimal sizes of code fragments (cf. Sect. 4.6). The clone detectors’ configurations are otherwise chosen in conformance with the literature; see also Table 2. All experiments have been performed on a Linux workstation with an Intel Core i7-10700K processor and 32 GB RAM.

### v2

@@ 2 Overall Approach and Architecture
Apparently, as can be seen by the above discussion, StoneDetector provides a rich set of configuration parameters that can be used for optimizing clone detection and demonstrates the tool’s versatility. Among the most critical parameters are the threshold value $\tau$ defining the required similarity of code clones, the minimum size $l$ of code fragments acceptable as code clones, the string metric $m$ used for path matching, and the hash function h used for description set encoding (cf. also Table 1). The therein given default values of configuration parameters are chosen for the tool’s best tradeoff of recall and precision with acceptable runtime. For an in-depth description of the StoneDetector’s configuration parameters, we refer the reader to the evaluation in Sect. 4 and to the supplementary information available online [^4] https://stonedetector.fmi.uni-jena.de/.
@@
String metric $\delta$ LCS LCSmodified
@@
Hashing algorithm $h$ none LSH
@@
Table 1: Main configuration parameters in StoneDetector
2.1 Extensibility of StoneDetector
@@
The use of fingerprints to represent description sets makes clone detection in StoneDetector fundamentally independent of the selected programming language and possible program representation. In fact, as is also depicted in Fig. 3, StoneDetector can be divided into a frontend and backend, with the former generating description sets of given code fragments and the latter performing the actual clone matching. The strict separation of frontend and backend allows us to facilitate a readily extensible tool, which can be easily extended for processing other programming languages as well as other internal program representations, such as control flow graphs, abstract syntax trees, or even just token sequences of linear code. Such an extension is only required to map the new programming language or program representations to fingerprints by using a certain hash function. Clone matching can subsequently be performed using the existing backend. The specific hash function used for construction is irrelevant to the functionality of the backend. In this way, fingerprints form a general and elegant interface for extending our system to other programming languages and program representations.
@@
To demonstrate the StoneDetector’s versatility and extensibility, we have added the handling of Java Bytecode as another frontend of StoneDetector [94] (cf. Fig. 3). The new frontend gets a set of class files as input and transforms them to fingerprint sets, which are committed to the existing clone matcher of StoneDetector. Specifically, processing class files is implemented via the Soot framework [107], which first transforms the contained Bytecode into the Bytecode-like Baf or into the register-based Jimple representation. Afterward, dominator trees are generated based upon the Baf or Jimple representation and paths contained in the dominator trees are extracted and encoded into fingerprint sets, as required for clone matching.
3 Clone Detection using Dominator Trees
3.1 Control flow and Dominator Trees
A control flow graph $G$ is a directed graph defined as $G=(N,E,s)$ with $(N,E)$ denoting the graph and $s\in N$ the start node [1]. We encode the instructions of a code fragment using the graph’s nodes. Note that we here use a single node for each instruction instead of using basic blocks. An edge $(k,l)\in E$ in a control flow graph connects two nodes $k$ and $l$ iff the instruction denoted by $l$ can be immediately executed after the instruction denoted by $k$. Each node $n\in N$ is then reachable via a path from the start node $s$, i.e., a sequence of consecutive edges $e\in E$ that begin at $s$ and end at $n$. An example of a control flow graph representing the lower code fragment shown in Fig. 2, i.e., function zipDir, can be found in Fig. 4 (a).
(a) Control Flow Graph
(b) Dominator Tree
(c) Abstract Dominator Tree
Dset= [[=, V, CALL, #0]->[=, V, NEWARRAY, byte]->[=, V, L]->[=, V, L]->[COND, LT, V, FIELDREAD]->
       [=, V, NEW, #1, V, V]->[COND, CALL, #2]->[CALL, #3, V, V, L],
       [=, V, CALL, #0]->[=, V, NEWARRAY, byte]->[=, V, L]->[=, V, L]->[COND, LT, V, FIELDREAD]->
       [=, V, NEW, #1, V, V]->[COND, CALL, #2]->[=, V, NEW,  #4, V]->[=, V, NEW, #5]->
       [CALL, #6, V]->[COND, NE, =, V, CALL, #7, V, NEG, V]->[CALL, #8, V, V, V],
       [=, V, CALL, #0]->[=, V, NEWARRAY, byte]->[=, V, L]->[=, V, L]->[COND, LT, V, FIELDREAD]->
       [=, V, NEW, #1, V, V]->[COND, CALL, #2]->[=, V, NEW,  #4, V]->[=, V, NEW, #5]->
       [CALL, #6, V]->[COND, NE, =, V, CALL, #7, V, NEG, V]->[CALL, #9],
       [=, V, CALL, #0]->[=, V, NEWARRAY, byte]->[=, V, L]->[=, V, L]->[COND, LT, V, FIELDREAD]->
       [=, V, NEW, #1, V, V]->[COND, CALL, #2]->[POSTINC, V]]
Fingerprints=
[35d37480aba8b098914b104b822d27af->d352f27168825f06297c0968af4006cc->c0812f2a5a758d7f6572a0f8fc057ce3->
 c0812f2a5a758d7f6572a0f8fc057ce3->20ae2bed041f26248c27148484197862->71bd8588d1df5ed91c8d27fb91a55d46->
 a45688ae3fc84165d19f2efe859e43db->c088b1671e63a009662f8e31c773a0a0,
 35d37480aba8b098914b104b822d27af->d352f27168825f06297c0968af4006cc->c0812f2a5a758d7f6572a0f8fc057ce3->
 c0812f2a5a758d7f6572a0f8fc057ce3->20ae2bed041f26248c27148484197862->71bd8588d1df5ed91c8d27fb91a55d46->
 a45688ae3fc84165d19f2efe859e43db->c21875dc52243cb4f19c52d3e02be40a->19b4a2a27bfa0ccbeeee5564571ec011->
 cb4b446eaed074420c4e9c663ffd3741->c15ffb85b338cca72f6729adcc304469,
 35d37480aba8b098914b104b822d27af->d352f27168825f06297c0968af4006cc->c0812f2a5a758d7f6572a0f8fc057ce3->
 c0812f2a5a758d7f6572a0f8fc057ce3->20ae2bed041f26248c27148484197862->71bd8588d1df5ed91c8d27fb91a55d46->
 a45688ae3fc84165d19f2efe859e43db->055bfda33fe11d9fb52471e283c27d9c->19b4a2a27bfa0ccbeeee5564571ec011->
 cb4b446eaed074420c4e9c663ffd3741->f13c29192ec234b0a51eb4f32bc85efe,
 35d37480aba8b098914b104b822d27af->d352f27168825f06297c0968af4006cc->c0812f2a5a758d7f6572a0f8fc057ce3->
 c0812f2a5a758d7f6572a0f8fc057ce3->20ae2bed041f26248c27148484197862->71bd8588d1df5ed91c8d27fb91a55d46->
 a45688ae3fc84165d19f2efe859e43db->fc6ebb514a9a13b92cff0ae97894c23e]
(d) Description Set and Fingerprints
Figure 4: Control flow graph, dominator tree, description set, and fingerprints of example
A dominator tree [1] is a tree and models a certain control flow property of code fragments, specifically, which instruction precedes the other under all possible executions of the respective code fragment. Thus, let the triple $G=(N,E,s)$ denote a certain control flow graph. A node $l\in N$ dominates a node $k\in N$ iff $l$ occurs on all paths from the start node $s$ to node $k$. Accordingly, every node $n\in N$ dominates itself by definition. A node $l\in N$ furthermore strictly dominates another node $k\in N$ iff $l$ dominates $k$ and $l\neq k$. Node $l$ is also immediately dominating $k$ iff $l$ strictly dominates $k$ and there exists no node $m\in N$, such that $m$ strictly dominates $k$ and $l$ strictly dominates $m$. The dominator tree for the control flow graph $G$ is then a directed graph $D=(N,E^{*})$ with $E^{*}=\{(l,k)\,|\,l\mbox{ immediately dominates }k\}$. An example of a dominator tree is visualized in Fig. 4 (b). As can be seen, the shown dominator tree corresponds to the control flow graph in Fig. 4(a). Note that in the dominator tree, each node is reachable from the start node, i.e., the tree’s root, via a single unambiguous path. This is in contrast to the given control flow graph. Also note that there are no backward edges encoding loop edges, and therefore cycles, in the dominator tree, since the loop body cannot dominate the loop condition according to the above definition.
We use an abstract, text-oriented encoding of instructions for our clone detection method, yielding the abstract dominator tree which is depicted in Fig. 4 (c). Therein, instructions and their respective operands are denoted in preorder notation. Furthermore, the distinct variable and field names, as well as literals, are not distinguished and mapped to a single joint symbol, respectively. Thus, all variable names are subsumed by symbol $V$, literals by symbol $L$, and so on, much like in the clone detector NiCad [85]. For example, consider instruction i = 0 which is encoded by [=,V,L] in the abstract dominator tree (cf. Fig. 4). Function calls are optionally represented using a constant pool. A constant pool access is denoted by the character # and the respective function’s entry in the constant pool, as, e.g., in in [CALL,#6,V] for instruction zos.putNextEntry(anEntry) (again, cf. Fig. 4). By using abstract dominator trees, we are able to more easily match code fragments that do not completely align with each other syntactically, in particular in the case of Type 2 and part of Type 3/4 code clone deviations.
3.2 Finding Similar Code Fragments
Our approach is similar to the method proposed by Alon et al. [6], where they use a linearization of abstract syntax trees based on paths that start and end at distinct leaf nodes of an abstract syntax tree. Though, in contrast, in our method, given a dominator tree $D=(N,E^{*})$, we consider all paths starting at the tree’s root node $s\in N$ and ending at a respective leaf node $l\in\{n\in N\,|\,\nexists(n,m)\in E^{*}\}$. The resulting set of these paths can then be used to unambiguously represent the dominator tree $D$. A single path is encoded by simply concatenating the (abstract) instructions denoted by the path’s nodes into a sequence of instructions. For example, considering the abstract dominator tree in Fig. 4 (c), the path from the start node to the rightmost leaf node is encoded by the sequence:
[…]
Let $D$ be a dominator tree, $\mathit{paths}_{D}$ the set of its paths, and $\mathit{leaves}_{D}$ the set of its leaf nodes in the following. The set of paths unambiguously representing $D$ is given by: $D_{\mathit{set}}=\{s\rightarrow\cdots\rightarrow n_{k}\,|\,s\cdots n_{k}\in\mathit{paths}_{D},n_{k}\in\mathit{leaves}_{D}\}$. We call the set $D_{\mathit{set}}$ the description set of dominator tree $D$. In the case of the abstract dominator tree depicted in Fig. 4 (c), we have four paths starting at the start node and ending in the leftmost, the left inner, the right inner, and the rightmost leaf node, respectively. Accordingly, the description set shown in Fig. 4 (d) contains the four instruction sequences representing the paths.
3.3 Metrics for Path Comparison
The Hamming distance [34] is a basic string metric, which is defined on text strings with fixed length. This distance measures the minimum number of substitutions to change one string into another and thus yields the number of deviating characters in two given strings. Therefore, as a limitation, the text strings are only compared up to the length of the respective shorter string, and the number of remaining characters in the longer string is merely added to the so far calculated distance measure. Note that the conventional Hamming distance therefore tends to penalize paths with different lengths. Note that the Hamming distance can be computed quite efficiently, which makes it a good candidate for fast and scalable near-miss code clone detection, as we have also observed in our experiments (cf. Sect. 4).
Example:
  • Code fragment 1: for (int j=0; j<<m; j++) { acc=acc-j; proc(acc); }
  • $f=\texttt{[=,V,L]}\rightarrow\texttt{[<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}\rightarrow\texttt{[=,V,+,V,L]}$
  • Code fragment 2: for (int l=0; l<<m; l++) { acc=acc+l; }
  • $g=\texttt{[=,V,L]}\rightarrow\texttt{[<,V,V]}\rightarrow\texttt{[=,V,+,V,V]}\rightarrow$ [=,V,+,V,L]
Hamming:
  • 1.
    Substitute [=,V,+,V,V] with [=,V,-,V,V]
    $g=\texttt{[=,V,L]}\rightarrow\texttt{[<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[=,V,+,V,L]}$
  • 2.
    Substitute [=,V,+,V,L] with [CALL,#4,V]
    $g=\texttt{[=,V,L]}\rightarrow\texttt{[<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}$
  • Hamming distance $\delta_{\mathit{H}}(f,g)=3=2+1$ because $f$ has one more character than $g$.
Levenshtein:
  • 1.
    Substitute [=,V,+,V,V] with [=,V,-,V,V]
    $g=\texttt{[=,V,L]}\rightarrow\texttt{[<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[=,V,+,V,L]}$
  • 2.
    Insert [CALL,#4,V]
  • $g=\texttt{[=,V,L]}\rightarrow\texttt{[<,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[\mbox{CALL},\#4,V]}\rightarrow\texttt{[=,V,+,V,L]}$
  • Levenshtein distance $\delta_{\mathit{L}}(f,g)=2$.
Needleman-Wunsch:
  • $f=\texttt{[=,V,L]}\rightarrow\texttt{[<,V,V]}\rightarrow\_\_\_\_\_\_\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}$ $\rightarrow\texttt{[=,V,+,V,L]}$
  • $g=\texttt{[=,V,L]}\rightarrow\texttt{[<,V,V]}\rightarrow\texttt{[=,V,+,V,V]}\rightarrow\_\_\_\_\_\_\rightarrow\_\_\_\_\_\_\rightarrow\texttt{[=,V,+,V,L]}$
  • Needleman-Wunsch-based distance $\delta_{\mathit{NW}}(f,g)=max(|f|-score(f,g),|g|-score(f,g))=2$, where $score(f,g)=3$.
Longest Common Subsequence (LCS):
  • $f=\texttt{[=,V,L]}\rightarrow\texttt{[\textless,V,V]}\rightarrow\texttt{[=,V,-,V,V]}\rightarrow\texttt{[CALL,\#4,V]}\rightarrow\texttt{[=,V,+,V,L]}$
  • $g=\texttt{[=,V,L]}\rightarrow\texttt{[\textless,V,V]}\rightarrow\texttt{[=,V,+,V,V]}\ \rightarrow\texttt{[=,V,+,V,L]}$
  • $z=\texttt{[=,V,L]}\rightarrow\texttt{[\textless,V,V]}\rightarrow\texttt{[=,V,+,V,L]}$
  • LCS-based distance $\delta_{\mathit{LCS}}(f,g)=max(|f|-|z|,|g|-|z|)=2$.
Figure 5: Distance metrics for calculating path similarity
Another metric that has its origins in the alignment of DNA sequences is the Needleman-Wunsch algorithm [71]. The algorithm computes an alignment of two given text strings that is optimal with respect to a certain scoring scheme using two steps. In the first step, a similarity matrix of the two text strings is generated based upon the scoring scheme, which defines weights for character matches, mismatches, and gaps. In the second step, the optimal string alignment with maximal total score is computed by performing a backward traversal on this similarity matrix. While the Levenshtein and Needleman-Wunsch-based distance require more computational effort, they are also more permissive in case of code clones with larger syntactical variance as in case of large-gap clones, where matching code fragments have many and/or larger edits similar to DNA sequence alignments [63].
The application of string metrics to the comparison of paths in abstract dominator trees is illustrated in Fig. 5 by way of an example. In the upper part, two code fragments are shown, conjoined with two paths $f$ and $g$ from their abstract dominator trees. As can be seen below, the (abstract) instructions take on the role of characters in the visualized calculations of string metrics. In the alignment provided for the Needleman-Wunsch algorithm, vertical lines between nodes represent matches and horizontal lines represent inserted gaps. For the value of $score(f,g)$, a scheme is used in which the weights for matches are set to $1$, for mismatches to 1-1, and for gaps to 0.
The Longest Common Subsequence (LCS) algorithm [12, 85] is a well-known tool for comparing the contents of files, used for example in the Unix diff utility or in the revision control system git. Besides, clone detectors like NiCad are relying on the LCS algorithm as well. In our case, the algorithm is applied to paths instead of lines of code, such that the longest common subpath is computed, e.g., $\texttt{[=,V,L]}\rightarrow\texttt{[<,V,V]}\rightarrow\texttt{[=,V,+,V,L]}$ for paths $f$ and $g$ in Fig. 5. This longest common subpath then forms the basis for defining the metric $\delta$ for two paths of dominator trees, which equals the maximum number of unique nodes for both paths.
Using the original LCS-based metric, though, can become problematic in cases of large-gap code clones [110], e.g., code clones where one code fragment is embedded into the other. The LCS metric then has problems recognizing the similarity of the respective paths of the dominator trees of the code fragments, as the paths differ too greatly in length. For this reason, a modified variant of the LCS-based metric can be used, which performs an additional similarity check in situations where two paths are not considered similar by the original LCS-based metric and therefore analyzes whether the length of the derived longest common subsequence is close to the length of the respective smaller path. If this is the case, the unique nodes in the smaller path are considered for building the LCS metric measure, rather than those of the larger path. Technically speaking, code fragments that consist of only one basic block are considered code clones if the calculated LCS value corresponds to a value that is less than a percentage of the number of nodes of the smaller code fragment that is specified by a certain threshold.
This additional detection feature for subclones can also be applied to such code fragments that do not consist of only one basic block. For that purpose, for two subpaths $q$ and $p$, for which the length of the longest common subsequence is close to the number of nodes of the smaller path $p$, the LCS value is determined using the uniquely occurring nodes in $p$. However, to avoid overly imprecise clone analysis, the examination for subpaths only makes sense if the length of the longest common subsequence of the two paths is nearby to the length of the smaller path $p$ and, furthermore, a minimum length is assumed for $p$. In our later described evaluation experiments (cf. Sect. 4.5), we observed that a maximum length difference of 10% in the number of nodes in the longest common subsequence and $p$ and a minimum length of 10 lines for $p$ is a good practical choice for these parameters.
Note that there exist other string metrics as well, besides the ones discussed above, which have also been implemented in StoneDetector. Among these are, e.g., N-gram distance, cosine distance, Jaccard similarity, etc. For the sake of brevity, we here limit ourselves to the subset of metrics that we use in the evaluation in Sect. 4. For a more comprehensive discussion of string metrics, we refer, e.g., to the survey [33].
3.4 Fingerprints and Hashing
However, using string metrics on the plain paths of (abstract) dominator trees for path comparison is rather inefficient due to the need for node matching, i.e., matching equal instructions, multiple times for a single comparison of two paths. Therefore, to minimize the costs, we conduct the node matching on fingerprints instead, which encode the respective nodes’ instructions. Such a fingerprint is simply the hashed value of an instruction, for which the string metrics are adjusted accordingly. There are apparently various hashing algorithms available for this purpose. For example, the MD5 algorithm [82] provides a widely used hash function with 16-byte hash values, which is often used as a checksum of file downloads and suffices for creating collision-free encodings of instructions in our abstract dominator trees. The MD5 algorithm is the default hashing algorithm for fingerprints in StoneDetector, and its application for encoding the description sets of the example in Fig. 4 is shown in the lower part of the same figure.
3.5 Comparison of Description Sets
Choosing one of the metrics in combination with a hashing algorithm, as presented in Sect. 3.3 and Sect. 3.4, for the distance function $\delta$ allows us to compare description sets and subsequently assess the similarity of code fragments. To this end, given two code fragments with respective description sets $D_{\mathit{set}}$ and $D^{\prime}_{\mathit{set}}$, the most similar path $q$ of $D^{\prime}_{\mathit{set}}$ is identified for each path $p$ of $D_{\mathit{set}}$, such that $q$ has a minimal value $\delta(p,q)$, and the arithmetic mean for these values over all paths $p\in D_{\mathit{set}}$ is computed. Formally, we can define the distance of the two description sets $D_{\mathit{set}},D^{\prime}_{\mathit{set}}\in\mathbb{D_{\mathit{set}}}$, having $|D_{\mathit{set}}|\leq|D^{\prime}_{\mathit{set}}|$ without loss of generality, as a function $\Delta$ such that:
@@ 4 Evaluation
In the evaluation, we are interested in finding answers to a number of research questions, in particular about the performance of our code clone detector compared to the current state of the art. We also present results that we believe to provide general insights into code clone detection for Java and to help other researchers when implementing code clone detectors, besides showing the versatility of StoneDetector. Thus, we address the following research questions:
@@
    Research Question 1: How many code clones does our clone detection method and tool StoneDetector detect, i.e., what is its recall compared to the state of the art, and how many of StoneDetector’s reported code clones are real code clones, i.e., what is the tool’s precision compared to the state of the art?
@@
    Research Question 2: Is the conventional code clone detector StoneDetector in general able to find more difficult-to-identify code clones, i.e., code clones with larger syntactical deviation?
@@
    Research Question 3: Does StoneDetector scale to large codebases containing up to multiple million lines of code?
@@
    Research Question 4: How do the used threshold value and the minimal size of code clone influence StoneDetector’s performance? Which role do different string metrics and hashing methods play?
@@
    Research Question 5: Can source code clones also be found in Java Bytecode using clone detectors like StoneDetector?
@@
CloneWorks 70% similarity threshold, aggressive configuration:
line splitting, blind renaming, literal abstraction
@@
In the comparison of StoneDetector with other code clone detectors for Java, a set of representative state-of-the-art tools has been chosen, including conventional text-based, token-based, tree and graph-based, as well as hybrid clone detectors (cf. Sect. 5). While we focus on conventional code clone detectors in this paper, we have also added the Oreo hybrid tool, which also implements a Deep Learning approach, for the sake of comparison. Namely, we use the following tools:
@@
In the evaluation experiments, we only consider code fragments with a minimal length of 15 lines in the original source code if not otherwise stated. Considering smaller code fragments in general results in a skewed comparison. For instance, in a Java method with only 6 lines of source code, half of the lines are usually just taken by the method’s annotations, signature, and curly braces. However, for the curious reader, we also provide experiments regarding different minimal sizes of code fragments (cf. Sect. 4.5). The clone detectors’ configurations are otherwise chosen in conformance with the literature; see also Table 2. In case of StoneDetector, the tool’ configuration has been chosen for the tool’s best tradeoff of recall and precision with acceptable runtime. When deviating from this default configuration, we will explain it in the respective sections. All experiments have been performed on a Linux workstation with an Intel Core i7-10700K processor and 32 GB RAM.